I am facing problem with all the question involving movable or multiple pulleys.Can anyone guide me how to face/approach such questions.


     

Grade:12

6 Answers

Shivam Dimri
11 years ago

best way to find the relation b/w accelarations of two bodies is to consider a constraint realtion b/w them,

or you can also consider the tension in the ropes which is directly realted to the accelerations of the bodies

if T is the tension pulling the pulley upwards (its attaced at the middle )

and T'' is the tension which is pulling the pulley downwards (the ropewhichis wound over the pulley)

then we get T  =  2T''

hence accelertions will also be a multiple of two!!!

if i was ableto make a diagram then it wyd be of great help...

Akash Kumar Dutta
11 years ago

apply constraint equations...refer DC Pandey for any help.

harihara kumar
11 years ago

the only way to solve pulley problems is by drawing free body diagram.ask ur lecturer how to draw a free body diagram.

Amar Bharti
11 years ago
yeah that is very tough to approach on own but with little help you''d be able to do almost all question : There is some basic "constraint relations" You can study the theory from any basic physics book ! It includes the understanding that if 1 block will go up how will the other (joined) blocks and things like that !
Amar Bharti
11 years ago
like in the condition (in the image): L+L''=constant If you''ll differentiate them Then V+V''=0 Differentiating again A+A''=0 You can use this result and apply newton''s law in any direction of motion Write the equation using the acceleration got from constraint relation :-)
vivek kumar
11 years ago

aplly correct constraint equation and connect the tensions in all strings with the help of fbd of pulley